the local convenience sells a large variety of snack foods. if their budget for the month include ls sales of $155,000 total variable cost of $77,500, and fixed costs of$23,500, how much must their sales be to broken even ?
Answer #1

Contribution margin =(155000-77500)/155000 =50%

Break-even point in units =Fixed costs/contribution margin

Break-even point in units =23500/50%

Break-even point in units =47000 units